# Unlocking Doping Effects on Altermagnetism in MnTe

### v1

@@ I Introduction
In this study, using density functional theory (DFT) calculations, we investigate how substitutional doping at the non-magnetic Te site influences the altermagnetic properties of MnTe. By designing supercells of MnTe, several doping configurations are created and thereby the DFT results are further analyzed in the context of the underlying symmetries of the doped MnTe systems.
Our results reveal that, although single-atom substitution lowers the system’s symmetry, it does not destroy altermagnetism-regardless of dopant type or position-due to the preserved six fold roto-inversion ($S_{6z}$) and mirror ($M_{z}$) symmetry, which maintains the $g$-wave altermagnetism. Pair doping leads to a plethora of possible configurations. Within the constraint of $2\times 2\times 2$ supercell considered here, there exist 120 possible configurations, which can be classified into five symmetry-based families. Among these, three families (46.66%) retain ideal altermagnetism, while the remaining two lack the necessary rotational or roto-inversion mirror symmetries connecting opposite spin sublattices and instead exhibit quasi-altermagnetic behavior. These quasi-altermagnetic states still show momentum-dependent spin splitting and nearly compensated magnetism. We identify the symmetry-breaking perturbations responsible for the transition from ideal to quasi-altermagnetic states using model Hamiltonian studies. Furthermore, pair-doped structures with lower symmetry exhibit anomalous Hall conductivity (AHC) when magnetization is oriented out-of-plane-a feature absent in pristine MnTe-while a quasi-altermagnetic configuration displays AHC along two different magnetization directions.
Overall, our study reveals how doping influences altermagnetism in MnTe, identifies the conditions and configurations required to preserve it or to break it, and highlights new opportunities for tuning AHC and altermagnetic properties through controlled chemical doping.
@@ II Structural and Computational Details
MnTe crystallizes in a hexagonal structure with alternating Mn and Te planes. Within each plane, Mn atoms are ferromagnetically coupled, while adjacent Mn planes are antiferromagnetically aligned along the $c$-axis, as illustrated in Fig. 1(a). The optimized lattice parameters for MnTe are $a$ = $b$ = 4.21 Å and $c$ = 6.70 Å, which closely match the experimentally reported values [43]. The $2\times 2\times 2$ supercell of MnTe was used to study effect of non-magnetic substitutional doping and is also shown in Fig.1(a).
@@ III.2 Magnetic structure of MnTe with single non-magnetic atom substitution
The local distortions significantly influence the electronic structure and magnetic properties. In pristine MnTe, the Mn atoms of the up (down) sublattice have LSM of 4.42 $\mu_{B}$ (-4.42 $\mu_{B})$ resulting in a zero net magnetization. In MnTe93.75Se6.25, Mn moments slightly vary from 4.42 to 4.45 $\mu_{B}$ due to the local perturbation induced by doping. However, we find that for a given LSM, there always exists an equal and opposite LSM. Such pairing [as shown in blue and red in Fig. 2(f)] gives rise to zero overall magnetization, as in the pristine case, which is further substantiated from the symmetric spin-resolved density of states (DOS) shown in Fig. 2(a), where the spin-up and spin-down DOS are symmetric at every energy. This naturally raises the point of possible observations of altermagnetism in the doped systems.
Figure 3: Altermagnetic and quasi-altermagnetic states in MnTe for different configurations with a pair of Se substitutions. Substituting a pair of Te atoms with Se atoms in a $\times2$$\times$2 MnTe supercell yields a system with 12.5% doping. This specific substitution results in 120 unique configurations, distributed among five distinct magnetic space groups (MSGs). The top panel [(a) and (b)] illustrate representative configurations from each MSG, categorized into two classes: (a) those exhibiting ideal altermagnetism, and (b) those displaying quasi-altermagnetic characteristics. The corresponding electronic band structures for these two classes are presented in panels (c) and (d), respectively. The valence and conduction band edges are shown in the insets, where the characteristic features of both altermagnetic and quasi-altermagnetic states are highlighted. Panel (e) provides a pie chart summarizing the distribution of the two-atom-doped configurations across the five MSGs. The gray, green, and yellow segments in the chart represent configurations that exhibit ideal altermagnetism, collectively accounting for approximately 46.66% of the total configurations.
@@
Figure 5: Anomalous Hall conductivity of pair doped MnTe with space group $Pm^{\prime}m^{\prime}2$ (No.25.60) and $C2^{\prime}/m^{\prime}2$ (No.12.62). Panels (a) and (b) show the anomalous Hall conductivity as a function of the Fermi level in pair-doped MnTe with the MSG $Pm^{\prime}m^{\prime}2$ (No. 25.60), for Sb and I dopants, respectively. The other two components, $\sigma_{yz}$ and $\sigma_{xz}$, vanish by symmetry. We note that the choice of dopant can effectively tune the magnitude of the anomalous Hall conductivity at the Fermi level. Panel (c) and (d) show the anomalous Hall conductivity for MSG $C2^{\prime}/m^{\prime}2$ (No.12.62) for Sb and I dopants, respectively. In this case, $\sigma_{yz}$ and $\sigma_{xy}$ are non-zero and $\sigma_{xz}$ vanishes. The change in pair-dopant position alters the active AHC component.
@@
In Fig. 2(c), we show the band structure of MnTe93.75Se6.25, along the altermagnetic $k$-path $\Gamma-LL^{\prime}$, where we observe spin-dependent splitting of the bands in the momentum space, characteristic of altermagnetism. For the visualization of the altermagnetic behavior in the full Brillouin zone, in Fig. 2(b), we have plotted the constant energy surface for the spin-up (blue) and spin-down (red) bands at $E=E_{F}-0.85$ eV. The coexistence of compensated magnetization and momentum-dependent spin-splitting confirms that MnTe retains its altermagnetic character even after substituting one Te atom with Se. To investigate whether the specific site of the Te substitution plays any role in determining this behavior, we performed additional calculations by substituting Se at different Te sites within the supercell. In all cases, the altermagnetic characteristics were preserved, indicating that the observed spin-splitting is intrinsic and independent of the substitution site.
@@
In a recent study, it has been shown that chemical bonding plays a major role in stabilizing and tuning the altermagnetism [25]. Therefore, we further explored the role of dopants by replacing Te with other non-isovalent atoms Sb and I. The respective band structures are shown in Fig. 2(d) and  2(e). Remarkably, in both the cases, we observe characteristic altermagnetic bands. Quantitatively, the substitution induces perturbations that lead to non-uniform magnetic moments among the Mn atoms within the supercell, similar to the Se atom doped case. For the Sb-doped case, Mn magnetic moments range from 4.42 to 4.45 $\mu_{B}$, and while for I doping, they range from 4.41 to 4.43 $\mu_{B}$. Despite small local variations, the spin-up and spin-down Mn sublattices remain magnetically compensated in both cases, similar to the case for Se doping, which implies the robustness of the altermagnetism. This invariance can be explained from a general symmetry perspective, as we illustrate next.
@@
The space group corresponding to the pristine MnTe is $P6_{3}/mmc$ (No.194), and corresponding MSG is $P6_{3}^{\prime}/m^{\prime}m^{\prime}c$ (No.194.268). Maintaining the same supercell geometry, the substitution of one Te atom in MnTe by another element reduces its symmetry, lowering the space group to $P\bar{6}m2$ (No.187). The corresponding MSG is $P\bar{6}^{\prime}m^{\prime}2$ (No.187.211). As a result of doping, the lowering of symmetry generates a larger number of symmetry-inequivalent Mn atoms. Since the Mn atoms exhibit varying LSM, the spin sublattice structure in the doped system is more complex, as shown in Fig. 2(f). Unlike the pristine case, where the up and down spin sublattices consist of only one Mn atom each, the doped system exhibits two spin sublattices, denoted spin-up in blue and spin-down in red, each containing 8 Mn atoms. These atoms are distributed across two different $c$-planes, with four atoms per plane. The spin-up and spin-down sublattices are related by a mirror symmetry operation ($M_{z}$) passing through the central plane, as illustrated in Fig. 2(f). The mirrored sites have equal and opposite LSM. Furthermore, the elements of spin-up and spin-down sublattices are connected through six-fold rotoinversion: a six-fold rotation about the $z$-axis followed by inversion through the dopant site. To give an example, in Fig. 2(f) and  2(g), we have illustrated how the rotoinversion takes Mn1 to Mn7, Mn2 to Mn5, Mn3 to Mn6, Mn4 to Mn8, and so on. We have further validated that the connecting sites have equal and opposite LSM, leading to a net zero magnetization. Therefore, from the symmetry perspective, this doped system is bound to exhibit altermagnetism. The formation of four nodal planes, three diagonals by rotoinversion ($S_{6z}$) and one basal (M001) by mirror symmetry leads to the emergence of of $g$-wave altermagnetism in the doped case.
@@
Next, we examine the electronic and magnetic structure of the doped system with a pair of dopants to gain further insights. While a single dopant resultant in one configuration irrespective of the position of the dopant in the supercell, a pair of dopants builds a large number of configurations defined by their relative positions. Moreover, a pair of dopants allows us to tailor the symmetry of both geometric and spin space groups, and therefore, we have a wider canvas to further explore the relationship between symmetry and altermagnetism. In the case of MnTe, a pair of dopants at non-magnetic sites gives us 120 possible configurations belonging to five magnetic space group families $C2^{\prime}/m^{\prime}$ (No.12.62), $Pm^{\prime}m^{\prime}2$ (No.25.60), $Fm^{\prime}m^{\prime}2$ (No.42.222), $P\bar{3}m^{\prime}1$ (No.164.89), and $P\bar{6}^{\prime}m^{\prime}2$ (No.187.211). In Fig. 3, one representative configurations from each of these five families and the distributions of the five families are shown. A further detailed symmetry analysis shows that, the configurations belongs to $P\bar{6}^{\prime}m^{\prime}2$ (No.187.211) has the ideal six-fold roto-inversion making them eligible to be ideal altermagnets, which is also directly validated from our calculated band structures shown in Fig. 3(c). A closer look at the structure reveals that this configuration has the same symmetry as of the single dopant case. However, we find two more families, namely $Pm^{\prime}m^{\prime}2$ (No.25.60), $Fm^{\prime}m^{\prime}2$ (No.42.222), where the band structures reflect all the characteristic of altermagnetism [see Fig. 3(a)]. Examining these structures further, we find that each of them can be represented through a smaller orthorhombic primitive unit cell with a single dopant in it. These unit cells exhibit two mirror symmetries connecting the opposite spin sublattices, which give rise to $d$-wave altermagnetism with two nodal planes in the reciprocal space. It is possible to make a transition from $d$-wave altermagnetism of the primitive unit cell to the $g$-wave altermagnetism of the supercell, considering that the latter is a hexagonal structure which carries three partially overlapped orthorhombic primitive unit cells (see Appendix- A). The two nodal planes from each of the three orthorhombic substructures lead to four distinct nodal planes, which result in $g$-wave altermagnetism.
@@
In the remaining two cases, corresponding to $C2^{\prime}/m^{\prime}$ (No.12.62) and $P\bar{3}m^{\prime}1$ (No.164.89), the situation differs from the three cases discussed earlier. Here, the relative positions of the dopants break the symmetry connecting opposite spin sublattices, as they are no longer related by rotation, translation, or inversion. Therefore, in an ideal context, they are no longer altermagnets. However, band structure analysis for the Se-doped case reveals an interesting observation [Fig. 3(d)]: the band structures exhibit a close resemblance to those of altermagnets, including spin-dependent splitting along the altermagnetic $k$-path, and even though the spin-splitting changes sign, it is not equal on either side of the nodal plane corresponding to the pristine system. We term these quasi-altermagnetic materials, augmenting the set of ideal altermagnets. These deviations are evident in these quasi-altermagnetic systems when comparing zoomed views of the valence band maximum and conduction band minimum near $\Gamma$ across all cases.
@@
Similar to the single-dopant case, we investigate the effects of hole and electron doping for the pair dopant cases. The band structures for the five MSGs doped with Sb (hole-doping) and I (electron-doping) are shown in Fig. 4. For MSGs that display altermagnetism in the isovalent-doping case, Sb and I doping also yield perfect altermagnetism. However, for the other MSGs, deviations from altermagnetic behavior are more pronounced with Sb and I. In the Sb-doped case, the conduction bands, as expected, show minimal deviation, whereas the valence bands exhibit significant changes. In the I-doped case, deviations occur in both valence band maximum and the conduction band minimum.
Although the ideal altermagnetic condition is broken for MSGs $C2^{\prime}/m^{\prime}$ (No.12.62) and $P\bar{3}m^{\prime}1$ (No.164.89), the net magnetic moment remains nearly zero for isovalent doping and very small for hole and electron doped cases. Band structure analysis along a random path containing both conventional and altermagnetic segments shows that the dominant spin channel alternates between spin-up and spin-down bands in different Brillouin zone regions, resulting in an overall moment close to zero.

### v2

@@ I Introduction
In this study, using density functional theory (DFT) calculations, we investigate how substitutional doping at the non-magnetic Te site influences the altermagnetic properties of MnTe. By designing supercells of MnTe, several doping configurations are created, and thereby the DFT results are further analysed in the context of the underlying symmetries of the doped MnTe systems.
Our results reveal that, although single-atom substitution lowers the system’s symmetry, it does not destroy altermagnetism-regardless of dopant type or position-due to the preserved six-fold roto-inversion ($S_{6z}$) and mirror ($M_{z}$) symmetry, which maintains the $g$-wave altermagnetism. Pair doping leads to a plethora of possible configurations. Within the constraint of $2\times 2\times 2$ supercell considered here, there exist 120 possible configurations, which can be classified into five symmetry-based families. Among these, three families (46.66%) retain ideal altermagnetism, while the remaining two lack the necessary rotational or roto-inversion mirror symmetries connecting opposite spin sublattices and instead exhibit quasi-altermagnetic behaviour. These quasi-altermagnetic states still show momentum-dependent spin splitting and nearly compensated magnetism. We identify the symmetry-breaking perturbations responsible for the transition from ideal to quasi-altermagnetic states using model Hamiltonian studies. Furthermore, pair-doped structures with lower symmetry exhibit anomalous Hall conductivity (AHC) when magnetization is oriented out-of-plane-a feature absent in pristine MnTe-while a quasi-altermagnetic configuration displays AHC along two different magnetization directions.
Overall, our study reveals how doping influences altermagnetism in MnTe, identifies the conditions and configurations required to preserve it or break it, reveals the emergence of a new magnetic subclass called quasi-altermagnets and its features, and highlights new opportunities for tuning AHC and altermagnetic properties through controlled chemical doping.
@@ II Structural and Computational Details
MnTe crystallizes in a hexagonal structure with alternating Mn and Te planes. Within each plane, Mn atoms are ferromagnetically coupled, while adjacent Mn planes are antiferromagnetically aligned along the $c$-axis, as illustrated in Fig. 1(a). The optimized lattice parameters for MnTe are $a$ = $b$ = 4.21 Å and $c$ = 6.70 Å, which closely match the experimentally reported values [43]. The $2\times 2\times 2$ supercell of MnTe was used to study the effect of non-magnetic substitutional doping and is also shown in Fig.1(a).
@@ III.2 Magnetic structure of MnTe with single non-magnetic atom substitution
Figure 3: Altermagnetic and quasi-altermagnetic states in MnTe for different configurations with a pair of Se substitutions. Substituting a pair of Te atoms with Se atoms in a $\times2$$\times$2 MnTe supercell yields a system with 12.5% doping. This specific substitution results in 120 unique configurations, distributed among five distinct MSGs. The top panel [(a) and (b)] illustrates representative configurations from each MSG, categorized into two classes:(a) those exhibiting ideal altermagnetism, and (b) those displaying quasi-altermagnetic characteristics. The corresponding electronic band structures for these two classes are presented in panels (c) and (d), respectively. The valence and conduction band edges are shown in the insets, where the characteristic features of both altermagnetic and quasi-altermagnetic states are highlighted. The constant energy surfaces for a pair of spin-up (blue) and spin-down channels (red) at $E=E_{F}-0.7$ eV for altermagnets and quasi-altermagnets are shown in panels (e) and (f), respectively. Constant energy surfaces of altermagnets illustrate the six-fold roto-inversion relation of opposite channels, whereas a distortion is visible in that of quasi-altermagnets. Panel (g) provides a pie chart summarizing the distribution of the two-atom-doped configurations across the five MSGs. The grey, green, and yellow segments in the pie chart represent configurations that exhibit ideal altermagnetism, collectively accounting for approximately 46.66% of the total configurations.
@@
The local distortions significantly influence the electronic structure and magnetic properties. In pristine MnTe, the Mn atoms of the up (down) sublattice have LSM of 4.42 $\mu_{B}$ (-4.42 $\mu_{B})$ resulting in a zero net magnetization. In MnTe0.9375Se0.6250, Mn moments slightly vary from 4.42 to 4.45 $\mu_{B}$ due to the local perturbation induced by doping. However, we find that for a given LSM, there always exists an equal and opposite LSM. Such pairing [as shown in blue and red in Fig. 2(f)] gives rise to zero overall magnetization, as in the pristine case, which is further substantiated from the symmetric spin-resolved density of states (DOS) shown in Fig. 2(a), where the spin-up and spin-down DOS are symmetric at every energy. This naturally raises the point of possible observations of altermagnetism in the doped systems.
@@
Figure 5: Schematic illustrations defining quasi-altermagnetism and its distinctness w.r.t. ferromagnetism, antiferromagnetism, and altermagnetism. The spin-split occurs in ferromagnetism due to a Zeeman like field. The rest of the configurations have sublattices, denoted here as A and B. The symmetry connections among these sublattices are expressed. The inversion symmetry in antiferromagnet introduces sub-band degeneracy, which is lifted in altermagnet due to the breakdown of the inversion symmetry. The equal but opposite momentum-dependent spin split mirroring a nodal plane in altermagnets is due to perfect rotational symmetries. Lack of this symmetry blurs the nodal plane and results in unequal splitting, which gives rise to quasi-altermagnetism. The $k$-dependent energy relations for each of the configurations are expressed. Here, $k_{\perp}$ denotes the perpendicular distance from the nodal plane.
Figure 6: Constant energy surfaces of altermagnet and quasi-altermagnet. Panels (a) and (b) show the constant energy surfaces of spin-up (blue) and spin-down (red) bands for pristine MnTe (MSG: $P6_{3}^{\prime}/m^{\prime}m^{\prime}c$ (No.194.268)) and Se-doped MnTe (MSG: $C2^{\prime}/m^{\prime}$ (No.12.62)), which is a quasi-altermagnet. The six-fold rotational symmetry connection between the opposite spin bands is clear for altermagnets, while its breaking is evident in the case of Se-doped MnTe, and these do not follow any well-defined symmetry relations.
@@
Figure 7: Anomalous Hall conductivity and Berry curvature of pair doped MnTe with space group $Pm^{\prime}m^{\prime}2$ (No.25.60) and $C2^{\prime}/m^{\prime}2$ (No.12.62). AHC as a function of the Fermi level for Sb-doped MnTe with MSG $Pm^{\prime}m^{\prime}2$ (No. 25.60) and $C2^{\prime}/m^{\prime}2$ (No. 12.62) in panels (a) and (b), and for I-doped cases in panels (c) and (d). Panels (e)–(h) show top views of Berry curvature components $\Omega_{x}$, $\Omega_{y}$, and $\Omega_{z}$ in the Brillouin zone. For Sb, berry curvature is plotted at $E=E_{F}-0.5$ eV, and for I it is at $E=E_{F}+0.15$ eV. For MSG 25.60, only $\sigma_{xy}$ is finite, while other components vanish; for MSG 12.62, $\sigma_{xy}$ and $\sigma_{yz}$ are non-zero and $\sigma_{xz}$ vanishes. Dopant type tunes the AHC magnitude at the Fermi level, while dopant pair positions control the active AHC tensor components.
@@
In Fig. 2(c), we show the band structure of MnTe0.9375Se0.6250, along the altermagnetic $k$-path $\Gamma-LL^{\prime}$, where we observe spin-dependent splitting of the bands in the momentum space, characteristic of altermagnetism. For the visualisation of the altermagnetic behaviour in the full Brillouin zone, in Fig. 2(b), we have plotted the constant energy surface for the spin-up (blue) and spin-down (red) bands at $E=E_{F}-0.85$ eV. The coexistence of compensated magnetization and momentum-dependent spin-splitting confirms that MnTe retains its altermagnetic character even after substituting one Te atom with Se. To investigate whether the specific site of the Te substitution plays any role in determining this behaviour, we performed additional calculations by substituting Se at different Te sites within the supercell. In all cases, the altermagnetic characteristics were preserved, indicating that the observed spin-splitting is intrinsic and independent of the substitution site.
@@
In a recent study, it has been shown that chemical bonding plays a major role in stabilizing and tuning the altermagnetism [25]. Therefore, we further explored the role of dopants by replacing Te with other non-isovalent atoms Sb and I. The respective band structures are shown in Fig. 2(d) and  2(e). Remarkably, in both the cases, we observe characteristic altermagnetic bands. Quantitatively, the substitution induces perturbations that lead to non-uniform magnetic moments among the Mn atoms within the supercell, similar to the Se atom doped case. For the Sb-doped case, Mn magnetic moments range from 4.42 to 4.45 $\mu_{B}$, while for I doping, they range from 4.41 to 4.43 $\mu_{B}$. Despite small local variations, the spin-up and spin-down Mn sublattices remain magnetically compensated in both cases, similar to the case for Se doping, which implies the robustness of the altermagnetism. This invariance can be explained from a general symmetry perspective, as we illustrate next.
@@
The space group corresponding to the pristine MnTe is $P6_{3}/mmc$ (No.194), and corresponding MSG is $P6_{3}^{\prime}/m^{\prime}m^{\prime}c$ (No.194.268). Maintaining the same supercell geometry, the substitution of one Te atom in MnTe by another element reduces its symmetry, lowering the space group to $P\bar{6}m2$ (No.187). The corresponding MSG is $P\bar{6}^{\prime}m^{\prime}2$ (No.187.211). As a result of doping, the lowering of symmetry generates a larger number of symmetry-inequivalent Mn atoms. Since Mn atoms exhibit varying LSM, the spin sublattice structure in the doped system is more complex, as shown in Fig. 2(f). Unlike the pristine case, where the up and down spin sublattices consist of only one Mn atom each, the doped system exhibits two spin sublattices, denoted spin-up in blue and spin-down in red, each containing 8 Mn atoms. These atoms are distributed across two different $c$-planes, with four atoms per plane. The spin-up and spin-down sublattices are related by a mirror symmetry operation ($M_{z}$) passing through the central plane, as illustrated in Fig. 2(f). The mirrored sites have equal and opposite LSM. Furthermore, the elements of spin-up and spin-down sublattices are connected through six-fold rotoinversion: a six-fold rotation about the $z$-axis followed by inversion through the dopant site. To give an example, in Fig. 2(f) and  2(g), we have illustrated how the rotoinversion takes Mn1 to Mn7, Mn2 to Mn5, Mn3 to Mn6, Mn4 to Mn8, and so on. We have further validated that the connecting sites have equal and opposite LSM, leading to a net zero magnetization. Therefore, from the symmetry perspective, this doped system is bound to exhibit altermagnetism. The formation of four nodal planes, three diagonals by rotoinversion ($S_{6z}$) and one basal (M001) by mirror symmetry leads to the emergence of $g$-wave altermagnetism in the doped case.
@@
Next, we examine the electronic and magnetic structure of the doped system with a pair of dopants to gain further insights. While a single dopant results in one configuration irrespective of the position of the dopant in the supercell, a pair of dopants builds a large number of configurations defined by their relative positions. Moreover, a pair of dopants allows us to tailor the symmetry of both geometric and spin space groups, and therefore, we have a wider canvas to further explore the relationship between symmetry and altermagnetism. In the case of MnTe, a pair of dopants at non-magnetic sites gives us 120 possible configurations belonging to five MSG families $C2^{\prime}/m^{\prime}$ (No.12.62), $Pm^{\prime}m^{\prime}2$ (No.25.60), $Fm^{\prime}m^{\prime}2$ (No.42.222), $P\bar{3}m^{\prime}1$ (No.164.89), and $P\bar{6}^{\prime}m^{\prime}2$ (No.187.211). In Fig. 3, one representative configurations from each of these five families and the distributions of the five families are shown. A further detailed symmetry analysis shows that, the configurations belonging to $P\bar{6}^{\prime}m^{\prime}2$ (No.187.211) have the ideal six-fold roto-inversion, making them eligible to be ideal altermagnets, which is also directly validated from our calculated band structures shown in Fig. 3(c). A closer look at the structure reveals that this configuration has the same symmetry as of the single dopant case. However, we find two more families, namely $Pm^{\prime}m^{\prime}2$ (No.25.60), $Fm^{\prime}m^{\prime}2$ (No.42.222), where the band structures reflect all the characteristic of altermagnetism [see Fig. 3(a)]. Examining these structures further, we find that each of them can be represented through a smaller orthorhombic primitive unit cell with a single dopant in it. These unit cells exhibit two mirror symmetries connecting the opposite spin sublattices, which give rise to $d$-wave altermagnetism with two nodal planes in reciprocal space. It is possible to make a transition from the $d$-wave altermagnetism of the primitive unit cell to the $g$-wave altermagnetism of the supercell, considering that the latter is a hexagonal structure which carries three partially overlapped orthorhombic primitive unit cells (see Appendix A). The two nodal planes from each of the three orthorhombic substructures lead to four distinct nodal planes, which result in $g$-wave altermagnetism.
@@
In the remaining two cases, corresponding to $C2^{\prime}/m^{\prime}$ (No.12.62) and $P\bar{3}m^{\prime}1$ (No.164.89), the situation differs from the three cases discussed earlier. Here, the relative positions of the dopants break the symmetry connecting opposite spin sublattices, as they are no longer related by rotation, translation, or inversion. Therefore, in an ideal context, they are no longer altermagnets. However, band structure analysis for the Se-doped case reveals an interesting observation [Fig. 3(d)]: the band structures exhibit a close resemblance to those of altermagnets, including spin-dependent splitting along the altermagnetic $k$-path, and even though the spin-splitting changes sign, it is not equal on either side of the nodal plane corresponding to the pristine system. These deviations are evident when comparing zoomed views of the valence band maximum and conduction band minimum near $\Gamma$ across all cases. We term these quasi-altermagnetic materials, augmenting the set of ideal altermagnets. Below in a separate subsection, we have comprehensively defined quasi-altermagnetism and explained how it is distinct from ferromagnetism, antiferromagnetism, and altermagnetism (see subsection III.4).
Similar to the band structure, the variations are also evident in the constant energy surface plots of quasi-altermagnets [Fig 3 (f)] compared to that of altermagnets [Fig 3 (e)]. Although the overall momentum-space spin texture of quasi-altermagnets exhibits a qualitative resemblance to the $g$-wave altermagnetic phase, notable differences appears. In the quasi-altermagnetic phase, the constant energy surface exhibits a slight imbalance, where either spin-up or spin-down bands dominate. Moreover, the perfectly matched pairs of spin-up and spin-down bands in altermagnets are absent in quasi-altermagnets.
@@
Similar to the single-dopant case, we investigate the effects of hole and electron doping for the pair dopant cases. The band structures for the five MSGs doped with Sb (hole-doping) and I (electron-doping) are shown in Fig. 4. For MSGs that display altermagnetism in the isovalent-doping case, Sb and I doping also yield perfect altermagnetism. However, for the other MSGs, deviations from altermagnetic behaviour are more pronounced with Sb and I. In the Sb-doped case, the conduction bands, as expected, show minimal deviation, whereas the valence bands exhibit significant changes. In the I-doped case, deviations occur in both the valence band maximum and the conduction band minimum.
So far, we have discussed the effect of symmetry lowering induced by non-magnetic atom substitutional doping on altermagnetic properties of MnTe. Substitutional doping at the magnetic atom site also leads to the symmetry lowering, and the resulting configurations exhibit altermagnetism or quasi-altermagnetism depending on the symmetry of the system determined by the relative position of dopant. A discussion regarding this is given in the Appendix  C. This primarily infers that if the roto-inversion is maintained with the doping at the magnetic site, the spin polarized defect bands also show altermagnetic characteristics.
III.4 Quasi-altermagnetism
Quasi-altermagnetism, as we define, refers to a magnetic phase that emerges when an altermagnet loses its rotational symmetries. This loss of symmetry can occur due to external stimuli such as strain, electric field, defects, and vacancies or chemical doping, as is the case in the present study. Like the altermagnets, quasi-altermagnets lack both translational symmetry and inversion symmetry. They do not exhibit Zeeman splitting as in the ferromagnets. Instead, they show an anisotropic momentum-dependent spin-splitting as in altermagnets. However, the splitting is not equal and opposite on either side of the nodal plane. In fact, the nodal plane itself is blurred and ill-defined in the case of quasi-altermagnets due to the fact that the sub-band degeneracy may get lifted on this plane. Furthermore, if the system is an insulator, the net magnetization remains zero, like the antiferromagnets and altermagnets. However, if it is metallic, it can have uncompensated magnetization. For example, MnTe1-xSex with MSG symmetry $C2^{\prime}/m^{\prime}2$ (No.12.62) and $P\bar{3}m^{\prime}1$ (No.164.89) are insulators which exhibit quasi-altermagnetism with zero net magnetization. However, MnTe1-xSbx and MnTe1-xIx with the same MSG symmetry are quasi-altermagnetic metals, and these compounds have uncompensated magnetization. Fig. 5 illustrates the key features of quasi-altermagnetism and how they differ from ferromagnetism, antiferromagnetism, and altermagnetism.
While, altermagnetism requires perfect rotational symmetries, quasi-altermagnetism is a perturbative deviation effect. The constant energy plots of different pairs of bands in altermagnetic and quasi-altermagnetic materials, shown in Fig. 6, illustrates that quasi-altermagnetism originates from the breakdown in rotational symmetry of altermagnets, rather that the superposition of different types of altermagnetic order in varied proportion. In most of the practical systems, it is therefore highly likely that quasi-altermagnetism exists instead of altermagnetism.